OpenStudy (anonymous):

2x^2+12x+19=(x+4)^2 Solve equation for y . (If there is more than one solution, separate them with commas.)

OpenStudy (noorik.):

what (y), the variable is (x). however, the right side (x+4)^2 = x^2 + 8x +16 then 2x^2+ 12x +19 = x^2 +8x +16 x^2 +4x +3 = 0 it's easy now, complete it yourself

OpenStudy (anonymous):

Okay, Thank you! (: